How many moles of calcium hydroxide will be produced if 8.76 miles of water reacts with calcium hydride?
1
Expert's answer
2020-03-11T09:06:34-0400

CaH2+2H2O=>Ca(OH)2+2H2

n(CaH2)=n(H2O)/2=n(Ca(OH)2)=4.38 moles